OpenStudy (anonymous):

Which of the following is the solution set of the given equation? 14 + 8m = 14 - 3m - 5m Ø {0} {all reals}

OpenStudy (john_es):

\[14+8m=14-3m-5m\Rightarrow 14+8m=14-8m\Rightarrow0=16m\Rightarrow m=0\]I

OpenStudy (anonymous):

thank you agaaainnn :))

OpenStudy (john_es):

;)